# Generalizable Trajectory Prediction via Inverse Reinforcement Learning with Mamba-Graph Architecture

### v1

@@ Abstract
Accurate driving behavior modeling is fundamental to safe and efficient trajectory prediction, yet remains challenging in complex traffic scenarios. This paper presents a novel Inverse Reinforcement Learning (IRL) framework that captures human-like decision-making by inferring diverse reward functions, enabling robust cross-scenario adaptability. The learned reward function is utilized to maximize the likelihood of output by the encoder-decoder architecture that combines Mamba blocks for efficient long-sequence dependency modeling with graph attention networks to encode spatial interactions among traffic agents. Comprehensive evaluations on urban intersections and roundabouts demonstrate that the proposed method not only outperforms various popular approaches in prediction accuracy but also achieves 2 times higher generalization performance to unseen scenarios compared to other IRL-based method.
@@ I Introduction
Modern intelligent transportation systems (ITS) rely on accurate trajectory prediction to enhance road safety and traffic efficiency [1]. However, predicting vehicle trajectories in complex urban environments remains a significant challenge due to the intricate interplay of factors, including dynamic road geometries, dense traffic interactions, and the inherent stochasticity of human driving behavior [2].
@@
Traditional prediction models with supervise learning paradigms aim to learn the mapping from historical trajectories to future trajectories, as well as the hidden state of the environment. Sequence modeling approaches, such as recurrent neural networks (RNNs) [3] and long short-term memory (LSTM) networks [4, 5, 6], have been widely employed to model temporal dependencies in trajectory data. With the advent of Mamba [7], a selective State Space Model (SSM), it offers superior capability in modeling long-range dependencies while maintaining exceptional inference efficiency—a critical attribute for real-world autonomous driving applications where computational latency is non-negligible. However, these methods often fall short in accurately capturing the intricate spatiotemporal interactions between vehicles and their environments. Graph-based encoding methods, such as Graph Neural Networks (GNNs) [8], have emerged as a promising solution to address the interrelationships between vehicles and their surroundings [9, 10, 11]. Yet, these models often struggle to capture the nuanced, context-dependent decision-making processes of human drivers, leading to suboptimal or unsafe predictions, particularly in scenarios involving complex maneuvers or unexpected events. Thus in this paper, we combined Mamba block and GNN in our predictor module.
@@
To further address the challenges of dataset dependency and cross-scenario generalization, Inverse Reinforcement Learning (IRL) [12] has shown promise in trajectory prediction complementing these supervised learning approaches. Whereas pure IRL approaches [13, 14], constrained by their reliance on environment dynamics modeling, are only viable in simplistic settings, our IRL-augmented data-driven approach not only extends applicability to complex environments, but also preserves high predictive accuracy. Diverging from conventional paradigms [15, 16]—which hinge on maximum margin theory and iteratively refine policies through learner-expert performance comparisons, thus viable solely for tasks with limited policy diversity—this work embraces the maximum entropy (MaxEnt) IRL to yield policies of maximal diversity, thereby conferring robustness against the noise and uncertainty endemic to human demonstrations, while simultaneously sustaining adaptability to dynamic driving scenarios (Fig. 1).
@@
  • Our method leverages the Mamba architecture to efficiently model long-range state-space dependencies in vehicle trajectories, augmented by GNN module that encodes inter-vehicle interactions through spatiotemporal message passing.
@@
The rest of paper is organized as follows: Section. II presents the proposed framework, detailing its core modules. Section. III describes the experimental setup, presents comprehensive validation results, and provides in-depth analysis of the findings. Finally, Section. IV concludes the paper with a summary of key contributions and discusses promising directions for future research.
@@ II Methodology
This section presents our proposed framework (Fig. 2), Environment-Aware Mamba Trajectory Predictor with MaxEnt IRL, which integrates sequence modeling with IRL to capture human driver behavior. We first formalize the trajectory prediction problem, then detail our encoder-decoder architecture leveraging Mamba blocks for efficient long-sequence modeling and GATs to encode spatial interactions. Finally, we introduce the MaxEnt IRL module to infer driver reward functions from heterogeneous human demonstrations, enabling more human-like and scenario-adaptive trajectory generation.
@@ II-A Problem Statement
Trajectory prediction is achieved by modeling and analyzing input historical trajectories, utilizing the model to forecast future trajectories. We employ $\tau$ to represent the historical trajectory of the target vehicle, while $\hat{\tau}$ denotes its future trajectory. Note that the analysis in this paper assumes a discrete framework, where a trajectory is defined as a sequence of states, i.e., $\tau=[s_{1},s_{2},\dots,s_{L}]$, with $s_{t}$ representing the state of the vehicle at the $t$-th time step such as position, velocity, yaw angle, acceleration. The discussion of trajectory representation will be extended to include inter-vehicle information expressed through graph-theoretic formalism in Section. II-B. $L$ indicates the length of the given trajectory.
While the use of reinforcement learning for prediction inherently involves uncertain approximations of the environment’s dynamic model, the historical trajectory input, which, even as time series data, implicitly encompasses state-action pairs, i.e., $a_{t}=\Delta x_{t}=x_{t+1}-x_{t}$ where $x_{t}$ is the coordinate position of the target vehicle contained in $s_{t}$. This permits the application of the Markov Decision Process (MDP) for modeling, and allows for the use of IRL to analyze the problem as detailed in Section. II-C.
Our goal is to predict the future trajectory of target vehicle, specifically the coordinate position $\{\hat{x_{t}}\}^{T}_{t=L}$, given historical trajectory $\tau\in\Gamma_{S}$, where $T$ is the final step of the prediction. Furthermore, we aims to implicitly infer a policy $\pi^{*}$ that maximizes the likelihood of generated trajectories $\hat{\tau}\sim\pi^{*}$,$\hat{\tau}\in\Gamma_{T}$, where $\Gamma_{S}$ and $\Gamma_{T}$ are the source and target domain of trajectory data. The potential discrepancy between source and target domains provides the fundamental motivation for investigating our method’s cross-scenario adaptability.
@@ II-B Graph-Based Mamba Predictor
Our prediction module adopts the widely utilized encoder-decoder architecture, wherein the embedded historical trajectory $\tau$ is processed by the encoder to yield a latent representation $h$ that encapsulates the spatiotemporal context, which subsequently serves as the input to the decoder for generating the trajectory $\hat{\tau}$. In the proposed architecture, the Mamba model is exclusively employed as the decoder, a deliberate design choice predicated on the observation that its formidable temporal representational capacity—while advantageous for sequential generation—would render an end-to-end fully Mamba-based encoder-decoder architecture prohibitively challenging to optimize.
The State Space Models map an input signal $x(t)\in\mathbb{R}^{D}$ to the output signal $y(t)\in\mathbb{R}^{D}$ via the latent state $h(t)\in\mathbb{R}^{N}$. For application to a discrete autoregressive input sequence of target vehicle states $(s_{0},s_{1},\dots)$ from each preceding timestep, our Mamba module iteratively updates its hidden state $h_{t}$ at each step $t$ (where $t$ ranges from $L+1$ to $T$):
@@
To complete the description of our encoder-decoder architecture, the encoder employs a GRU network while following the methodology of [17, 11] to integrate inter-vehicle information into the latent representation $h_{L}$ through a Graph Attention Network (GAT)[18]. Inter-vehicle relationships are modeled as graph $\mathcal{G}=(\mathcal{V},\mathcal{E})$, where $\mathcal{V}$ denotes the node set, $\mathcal{E}$ denotes the edge set. Each node represents a traffic participant. For feature representation, let $h^{\nu}\in\mathbb{R}^{d_{f}}$ be the features of node $\nu\in\mathcal{V}$, $e^{\nu,\iota}\in\mathbb{R}^{d_{e}}$ be the features of edge $(\nu,\iota)\in\mathcal{E}$. This design choice is particularly crucial as the relational information between target and surrounding vehicles is typically provided as the node and edge features in the heterogeneous graph.
Within this framework, we implement an attention mechanism that computes aggregation weights across each node $\nu$’s inclusive neighborhood $\tilde{\mathcal{N}}(\nu)=\mathcal{N}(\nu)\cup\{\nu\}$, enabling the model to allocate varying importance to different neighbors in the graph. Specifically, we utilize an enhanced attention computation where the attention weights are determined by:
@@
where $\mathbf{a}_{\alpha}$ and $\mathbf{W}_{\alpha}$ are learnable parameters of GATs; $\gamma(\cdot)$ denotes the Leaky ReLU activation function (with negative slope $0.2$), providing non-linear transformation; $[h^{\nu}\|h^{\iota}][e_{\nu,\iota}]$ indicates concatenation of features. The hidden state of the target vehicle node with graph considered is updated through:
@@ II-C Mamba-based MaxEnt IRL
As delineated in Section. II-A, given $N$ observed trajectories $\Gamma=\{\tau_{i}\}^{N}=\{s_{1},s_{2},\dots,s_{L}\}_{N}$, which can be transformed into a set of human demonstrations $\mathcal{D}=\{(s_{t},a_{t})\}^{L\times N}_{t=1}$, our principal objective revolves around the estimation of the underlying reward function, the accurate reconstruction of which would enable the emulation and prediction of nuanced human driving behaviors with enhanced fidelity[19].
@@
where $\pi^{*}$ is the human strategy within the human demonstration. The trajectory likelihood $P(\tau)$, governed by the maximum entropy principle, is proportional to the exponential of the trajectory’s cumulative reward. To elaborate with precision, the likelihood of a trajectory $\tau$ necessitates normalization through division by the partition function $Z(\theta_{\text{RF}})$—a quantity explicitly defined in Eq. (6). Here with a bit abuse of notation, the summation $\sum_{\tau}$ enumerates over the entire space of feasible human trajectories, while $R(\tau|\theta_{\text{RF}})$ encapsulates the cumulative rewards associated with a specific trajectory $\tau$ parameterized by $\theta_{\text{RF}}$. This normalization not only guarantees the requisite properties of a probability measure but also introduces a logarithmic regularization term during optimization, thereby mitigating the risk of overfitting to the human demonstrations.
@@
In practical applications, computing the partition function $Z$ is generally infeasible since the trajectory space is typically extremely large. To circumvent this computational impasse, one may resort to importance sampling [19] techniques, whereby $Z$ is approximated through Monte Carlo estimation using $M$ samples from $\Gamma$:
@@
The Trajectory Prediction Module (TPM) integrates all components described in Section. II-B, namely the graph-based encoder and Mamba decoder. The TPM updates its parameters by simultaneously optimizing the mean squared error of the predicted coordinates, while also maximizing the cumulative reward of the predicted trajectory in accordance Eq. (5):
@@
6:     Compute reward values from trajectores for maximum likelihood in both $\mathcal{L}_{\text{TPM}}$ and $\mathcal{L}_{\text{RF}}$:  - Historical trajectory rewards $\{r_{i}\}_{t=1}^{L}=R(\tau|\theta_{RF})$  - Predicted trajectory rewards $\{r_{i}\}_{t=L}^{T}=R(\hat{\tau}|\theta_{RF})$

### v2

@@ Abstract
Accurate driving behavior modeling is fundamental to safe and efficient trajectory prediction, yet remains challenging in complex traffic scenarios. This paper presents a novel Inverse Reinforcement Learning (IRL) framework that captures human-like decision-making by inferring diverse reward functions, enabling robust cross-scenario adaptability. The learned reward function is utilized to maximize the likelihood of output by integrating Mamba blocks for efficient long-sequence dependency modeling with graph attention networks to encode spatial interactions among traffic agents. Comprehensive evaluations on urban intersections and roundabouts demonstrate that the proposed method not only outperforms various popular approaches in terms of prediction accuracy but also achieves 2.3 times higher generalization performance to unseen scenarios compared to other baselines, achieving adaptability in Out-of-Distribution settings that is competitive with fine-tuning.
@@ I Introduction
Modern intelligent transportation systems (ITS) rely on accurate trajectory prediction to enhance road safety and traffic efficiency [20]. However, predicting vehicle trajectories in complex urban environments remains a significant challenge due to the intricate interplay of factors, including dynamic road geometries, dense traffic interactions, and the inherent stochasticity of human driving behavior [21]. Beyond these challenges, deploying trajectory prediction in autonomous driving requires addressing the need for fast and accurate trajectory prediction in unseen, even Out-of-Distribution (OOD) [19] scenarios. This necessitates handle the domain shift between training dataset and real-world deployment conditions.
@@
Traditional prediction models with supervise learning paradigms aim to learn the mapping from historical trajectories to future trajectories, as well as the hidden state of the environment. Sequence modeling approaches, such as recurrent neural networks (RNNs) [25] and long short-term memory (LSTM) networks [17], have been widely employed to model temporal dependencies in trajectory data. With the advent of Mamba [8], a selective State Space Model (SSM), it offers superior capability in modeling long-range dependencies while maintaining exceptional inference efficiency—a critical attribute for real-world autonomous driving applications where computational latency is non-negligible. However, these methods often fall short in accurately capturing the intricate spatio-temporal interactions between vehicles and their environments. Graph-based encoding methods, such as Graph Neural Networks (GNNs) have emerged as a promising solution to address the interrelationships between vehicles and their surroundings [22]. Yet, these models often struggle to capture the nuanced, context-dependent decision-making processes of human drivers, leading to suboptimal or unsafe predictions, particularly in scenarios involving complex maneuvers or unexpected events. Thus in this paper, we combined Mamba block and GNN in our predictor module.
@@
To further address the challenges of dataset dependency and cross-scenario generalization, Inverse Reinforcement Learning (IRL) has shown promise in trajectory prediction complementing these supervised learning approaches. Whereas pure IRL approaches [10], constrained by their reliance on environment dynamics modeling, are only viable in simplistic settings, our IRL-augmented data-driven approach not only extends applicability to complex environments, but also preserves high predictive accuracy. Diverging from conventional paradigms [7]—which hinge on maximum margin theory and iteratively refine policies through learner-expert performance comparisons, thus viable solely for tasks with limited policy diversity—this work embraces the maximum entropy (MaxEnt) IRL to yield policies of maximal diversity, thereby conferring robustness against the noise and uncertainty endemic to human demonstrations, while simultaneously sustaining adaptability to dynamic driving scenarios.
In OOD scenarios, where the environmental data distribution completely differs from the training dataset, most data-driven methods experience significant performance degradation. Furthermore, recent studies have highlighted the Mamba architecture’s limitations in generalization, such as constrained length extrapolation [2] and limited domain generalization capabilities in visual tasks [14]. Despite these challenges, by leveraging the human driving reward function acquired through IRL, our approach integrates an explicit policy extension using established off-policy Reinforcement Learning (RL) algorithm. This integration substantially improves trajectory prediction performance in OOD settings, achieving results comparable to the fine-tuned method that utilize ground truth data from new environments—all without requiring any actual ground truth data from the target scenarios (Fig. 1).
@@
  • Our method leverages the Mamba architecture to efficiently model long-range state-space dependencies in vehicle trajectories, augmented by GNN module that encodes inter-vehicle interactions through spatio-temporal message passing.
@@
  • We extend the generalization capability of our method by incorporating an explicit policy submodule, which achieves OOD adaptability comparable to fine-tuning approaches, all without substantial computational overhead.
The rest of paper is organized as follows: Section II presents the proposed framework, detailing its core modules. Section III describes the experimental setup, presents validation results, and provides in-depth analysis of the findings. Section IV concludes the paper with a summary of key contributions and discusses promising directions for future research.
@@ II Methodology
This section presents our proposed framework (Fig. 2), Environment-Aware Mamba Trajectory Predictor with MaxEnt IRL, which integrates sequence modeling with IRL to capture human driver behavior. We first formalize the trajectory prediction problem, then detail our encoder-decoder architecture leveraging Mamba blocks for efficient long-sequence modeling and Graph Attention Network (GAT) to encode spatial interactions. Finally, we introduce the MaxEnt IRL module to infer driver reward functions from heterogeneous human demonstrations, enabling more human-like and scenario-adaptive trajectory generation.
@@ II-A Problem Statement
We employ $\tau$ to represent the historical trajectory of the target vehicle, while $\hat{\tau}$ denotes its future trajectory. Note that the analysis in this paper assumes a discrete framework, where a trajectory is defined as a sequence of states, i.e., $\tau=[s_{1},s_{2},\dots,s_{L}]$, with $L$ indicating the length of the given trajectory and $s_{t}$ representing the state of the vehicle at the $t$-th time step such as position, velocity, yaw angle, acceleration. The discussion of trajectory representation will be extended to include inter-vehicle information expressed through graph-theoretic formalism in Section II-B.
While the RL-based prediction involves uncertain approximations of the environmental dynamics, the historical trajectory input, which, even as time series data, implicitly encompasses state-action pairs, i.e., $a_{t}=\Delta x_{t}=x_{t+1}-x_{t}$ where $x_{t}$ is the coordinate position of the target vehicle contained in $s_{t}$. This treatment abstracts the decision-making to a behavioral level, capturing high-level intent and social interactions rather than low-level control signals. Under high-frequency sampling, displacement corresponds to instantaneous velocity, and we implicitly assume a near-ideal underlying controller—making the mapping from intent $a_{t}$ to state change $s_{t+1}-s_{t}$ effectively deterministic. Thus, geometric displacement is justified as a valid “action,” enabling the use of an MDP formulation and IRL analysis as detailed in Section II-C.
Our goal is to predict the future trajectory of target vehicle, specifically the coordinate position $\{\hat{x_{t}}\}^{T}_{t=L}$, given historical trajectory $\tau\in\Gamma_{S}$, where $T$ is the final step of the prediction. Furthermore, we aims to implicitly infer a policy $\pi^{*}$ that maximizes the likelihood of generated trajectories $\hat{\tau}\sim\pi^{*}$$\hat{\tau}\in\Gamma_{T}$, where $\Gamma_{S}$ and $\Gamma_{T}$ are the source and target domain of trajectory data. The potential discrepancy between source and target domains provides the fundamental motivation for investigating our method’s cross-scenario adaptability.
@@ II-B Graph-Based Mamba Predictor
Our prediction module adopts the widely utilized encoder-decoder architecture, wherein the embedded historical trajectory $\tau$ is processed by the encoder to yield a latent representation $h$ that encapsulates the spatio-temporal context, which subsequently serves as the input to the decoder for generating the trajectory $\hat{\tau}$. In the proposed architecture, the Mamba model is exclusively employed as the decoder, a deliberate design choice predicated on the observation that its formidable temporal representational capacity—while advantageous for sequential generation—would render an end-to-end fully Mamba-based encoder-decoder architecture prohibitively challenging to optimize.
The State Space Models map an input signal $u(t)\in\mathbb{R}^{D}$ to the output signal $y(t)\in\mathbb{R}^{D}$ via the latent state $h(t)\in\mathbb{R}^{N}$. For application to a discrete autoregressive input sequence of target vehicle states $(s_{0},s_{1},\dots)$ from each preceding timestep, our Mamba module iteratively updates its hidden state $h_{t}$ at each step $t$ (where $t$ ranges from $L+1$ to $T$):
@@
To complete the description of our encoder-decoder architecture, the encoder employs a Gate Recurrent Unit (GRU) to integrate inter-vehicle information into the top of GRU output-latent representation $h_{L}$, through a GAT. Inter-vehicle relationships are modeled as graph $\mathcal{G}=(\mathcal{V},\mathcal{E})$, where $\mathcal{V}$ denotes the node set, $\mathcal{E}$ denotes the edge set. Each node represents a traffic participant. For feature representation, let $h^{\nu}\in\mathbb{R}^{d_{f}}$ be the features of node $\nu\in\mathcal{V}$, $e^{\nu,\iota}\in\mathbb{R}^{d_{e}}$ be the features of edge $(\nu,\iota)\in\mathcal{E}$. This design choice is particularly crucial as the relational information between target and surrounding vehicles is typically provided as the node and edge features in the heterogeneous graph.
Within this framework, we implement an attention mechanism that computes aggregation weights across each node $\nu$’s inclusive neighborhood $\tilde{\mathcal{N}}(\nu)=\mathcal{N}(\nu)\cup\{\nu\}$, enabling the model to allocate varying importance to different neighbors. Specifically, we utilize an enhanced attention computation where the attention weights $\alpha_{\nu,\iota}$ are determined by:
@@
where $\mathbf{a}_{\alpha}$ and $\mathbf{W}_{\alpha}$ are learnable parameters of GATs; $\gamma(\cdot)$ denotes the Leaky ReLU activation function (with negative slope $0.2$), providing non-linear transformation; $h^{\nu}\|h^{\iota}][e_{\nu,\iota}]$ indicates concatenation of features. With graph considered, the output of GRU, i.e., hidden state $h^{\prime\nu}$ of the target vehicle node $\nu$, is subsequently updated through:
@@ II-C Mamba-based MaxEnt IRL
As delineated in Section II-A, given $N$ observed trajectories $\Gamma=\{\tau_{i}\}^{N}=\{s_{1},s_{2},\dots,s_{L}\}_{N}$, which can be transformed into a set of human demonstrations $\mathcal{D}=\{(s_{t},a_{t})\}^{L\times N}_{t=1}$, our principal objective revolves around the estimation of the underlying reward function, the accurate reconstruction of which would enable the emulation and prediction of nuanced human driving behaviors with enhanced fidelity.
@@
where $\pi^{*}$ is the human strategy within the human demonstration. The trajectory likelihood $P(\tau)$, governed by the maximum entropy principle, is proportional to the exponential of the trajectory’s cumulative reward. To elaborate with precision, the likelihood of a trajectory $\tau$ necessitates normalization through division by the partition function $Z(\theta_{\text{RF}})$ as defined in Eq. (7). Here with a bit abuse of notation, the summation $\sum_{\tau}$ enumerates over the entire space of feasible human trajectories, while $R(\tau|\theta_{\text{RF}})$ encapsulates the cumulative rewards associated with a specific trajectory $\tau$ parameterized by $\theta_{\text{RF}}$. This normalization not only guarantees the requisite properties of a probability measure but also introduces a logarithmic regularization term during optimization, thereby mitigating the risk of overfitting to the human demonstrations.
@@
Directly computing the partition function $Z$ over the vast trajectory space is typically infeasible. To circumvent this, we approximate $Z$ via importance sampling using Monte Carlo estimation with $M$ samples from $\Gamma$:
@@
The Trajectory Prediction Module (TPM) integrates all components described in Section II-B, namely the graph-based encoder and Mamba decoder. The TPM updates its parameters by simultaneously optimizing the mean squared error of the predicted coordinates, while also maximizing the cumulative reward of the predicted trajectory in accordance Eq. (5):
@@
6:  Compute reward values from trajectories for maximum likelihood in both $\mathcal{L}_{\text{TPM}}$ and $\mathcal{L}_{\text{RF}}$:  - Historical trajectory rewards $\{r_{i}\}_{t=1}^{L}=R(\tau|\theta_{RF})$  - Predicted trajectory rewards $\{r_{i}\}_{t=L}^{T}=R(\hat{\tau}|\theta_{RF})$
@@
II-D Out-of-Distribution Generalization For Mamba Predictor
In OOD scenarios, directly predicting accurate future vehicle trajectory coordinates becomes particularly challenging. To achieve broader scenario coverage, we introduce an explicit policy extension, by employing the learned driver policy to produce sequential decisions over future states to generate predicted trajectories. Specifically, the driver agent takes the predicted future states as input, outputs predicted actions $a=\pi_{\theta_{\text{agent}}}(s)$ which align with human driving strategies, and subsequently produces predicted future vehicle coordinates. A detailed description of this process is provided in Algorithm. 2. Although the cumulative offset operation in the diagram is a for-loop, parallel computation is implemented in trajectory level and batch level, significantly reducing processing time. Parameter $\theta_{\text{agent}}$ is optimized using the off-policy Actor-Critic algorithm Twin Delayed Deep Deterministic Policy Gradient (TD3) [6].
To support this learning process, we construct an experience replay buffer with reward value labeled for each state-action pairs in vehicle trajectories using the reward function derived earlier. We further introduce a penalty term in the reward proportionally to their prediction errors (e.g., negative Mean Squared Error). Concretely, for each state-action pair with state $s_{t}$ and action $a_{t}$, we complete the transition with reward:
$r=R(s_{t},a_{t}|\theta_{\text{RF}})-\left\|a_{t}-\pi_{\theta_{\text{agent}}}(s_{t})\right\|^{2}$ (10)
Algorithm 2 Predict Trajectories in OOD Scenario with Explicit Policy
0: Historical OOD trajectory $\tau=[s_{1},s_{2},\dots,s_{L}]$ (each $s_{t}$ contains coordinate $x_{t}$), Optimized $\theta_{\text{agent}}^{*}$, $\theta_{\text{TPM}}^{*}$
0: Future coordinates $\{\hat{x}_{t}\}_{t=L+1}^{L+T}$
1: Predict future states: $s_{\text{future}}=[s_{L+1},s_{L+2},\dots,s_{L+T}]\leftarrow\mathcal{P}_{\theta_{\text{TPM}}^{*}}(\tau)$
2: Initialize action sequence $A\leftarrow\emptyset$
3:for each predicted state $s_{t}\in s_{\text{future}}$ do
4:  Generate action: $a_{t}\leftarrow\pi_{\theta_{\text{agent}}^{*}}(s_{t})$
5:  Append to action sequence: $A\leftarrow A\cup\{a_{t}\}$
6:end for
7: Initialize future coordinates $\{\hat{x}_{t}\}_{t=L+1}^{L+T}\leftarrow\emptyset$ and cumulative offset $\delta\leftarrow 0$
8: Let $x_{L}$ be the coordinate from the last historical state $s_{L}$
9:for each action $a_{t}\in A$ with index $t$ from $L+1$ to $L+T$ do
10:  Update cumulative offset: $\delta\leftarrow\delta+a_{t}$
11:  Compute future coordinate: $\hat{x}_{t}\leftarrow x_{L}+\delta$
12:  Add to output: $\{\hat{x}_{t}\}_{t=L+1}^{L+T}\leftarrow\{\hat{x}_{t}\}_{t=L+1}^{L+T}\cup\{\hat{x}_{t}\}$
13:end for
14:return $\{\hat{x}_{t}\}_{t=L+1}^{L+T}$